OpenStudy (anonymous):

Algebra 2 Heeelllpppp!!!! A rental car agency charges $14 per day plus 12 cents per mile to rent a certain car. Another agency charges $20 per day plus 8 cents per mile to rent the same car. How many miles per day will have to be driven for the cost of a car from the first agency to equal the cost of a car from the second agency? This is a review for a test so i really need to understand how to do this.

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

first agency : charges $ 14 per day + 12 cents per mile (m) ---- 14 + 0.12m 2nd agency : charges $ 20 per day + 8 cents per mile ---- 20 + 0.8m now we set them equal and solve for m to find out how many miles will make the car agencies charge the same. 14 + 0.12m = 20 + 0.08m -- subtract 14 from both sides 0.12m = 20 - 14 + 0.08m -- subtract 0.08m from both sides 0.12m - 0.08m = 20 - 14 -- combine like terms .04m = 6 -- divide both sides by .04 m = 150 check.. 14 + 0.12m = 20 + 0.08m 14 + 0.12(150) = 20 + 0.08(150) 14 + 18 = 20 + 12 32 = 32 (correct) Both agencies will cost the same when 150 miles are driven